a particle of mass m, intially at rest at origin is acted upon by a force F=kx^2. Find the speed of particle when it is at pont(L,0,0)
 

Grade:

1 Answers

Abhishek Kumar
9 years ago
Since,
Work done by all forces = change in KE
Co, change in KE = = \int kx^{2}.dx = kx^{3}/3 = kL3/3
Since initial KE is zero , Final KE = Change in KE = kL3/3